OpenStudy (anonymous):

show how y^2=2+xy turns into dy/dx=y/(2y-x)

OpenStudy (anonymous):

By differentiating implicitly 2y(dy/dx)=y+x(dy/dx) (dy/dx)(2y-x)=y dy/dx=y/(2y-x)